Use the given conditions to write an equation for the line in​ point-slope form and in​ slope-intercept form.
Passing through (6,6) with x-intercept 4
Write an equation for the line in point-slope form:
Write an equation for the line in slope-intercept form:
Thank you <3

Answer by stanbon(75887)

Use the given conditions to write an equation for the line in​ point-slope form and in​ slope-intercept form.
Passing through (6,6) with x-intercept (4,0)
Write an equation for the line in point-slope form:
y-6 = slope(x-6) where slope = (6-0)/(6-4) = 3
y-6 = 3x-18
y = 3x - 12
-------------------
Write an equation for the line in slope-intercept form:
slope = 3
Form:: y = mx + b
Solve for "b"::
6 = 3*6 + b
b = -12
y = 3x - 12
---------------
Cheers,
Stan H.
-------------